# Dynamical tide modified Roche limit in eccentric, asynchronous binaries

### v1

@@ Appendix B Surface terms in the third order perturbations of the gravitational potential energy
At the formulation level, it is clear that the affine model agrees with that in T. Van Hoolst (1994) when applied to incompressible stars. However, it is not trivial to show the agreement, especially for terms of $\mathcal{O}(\xi^{3})$ and above. This is because these terms in T. Van Hoolst (1994) often involve high-order derivatives of the quantities from the background or the lower perturbation levels, which produce surface terms that require integrating the Dirac delta functions. In this appendix, we explicitly calculate the $\mathcal{O}(\xi^{3})$ terms in the gravitational potential energy and show their agreement with the affine model under careful treatment of the surface terms.
@@
The Lagrangian displacement expanded in terms of the quadrupolar f-mode eigenfunctions is given by Eq. (7). The integral of three spherical harmonics can be evaluated using the Wigner 3-j symbols (see, e.g., N. N. Weinberg et al. (2012)):

### v2

@@ Appendix B Surface terms in the third order perturbations of the gravitational potential energy
At the formulation level, it is clear that the affine model agrees with that in Van Hoolst (1994) when applied to incompressible stars. However, it is not trivial to show the agreement, especially for terms of $\mathcal{O}(\xi^{3})$ and above. This is because these terms in Van Hoolst (1994) often involve high-order derivatives of the quantities from the background or the lower perturbation levels, which produce surface terms that require integrating the Dirac delta functions. In this appendix, we explicitly calculate the $\mathcal{O}(\xi^{3})$ terms in the gravitational potential energy and show their agreement with the affine model under careful treatment of the surface terms.
@@
The Lagrangian displacement expanded in terms of the quadrupolar f-mode eigenfunctions is given by Eq. (8). The integral of three spherical harmonics can be evaluated using the Wigner 3-j symbols (see, e.g., Weinberg et al. (2012)):